OpenStudy (anonymous):

A store mixes Brazilian coffee worth $10 per kilogram and Turkish coffee worth $12 per kilogram. The mixture is to sell for $11 per kilogram. Find how much of each should be used to make a 360-kilogram mixture How many kg of Brazilian coffee should be in the mixture? How many kg of Turkish coffee should be in the mixture?

OpenStudy (anonymous):

180 180 not sure

OpenStudy (anonymous):

do you mean the mixture is worth 11 per

OpenStudy (anonymous):

I assume you want a mixture that is worth $11/kg. Since Brazilian coffee is worth 10, and Turkish is worth 12, 11 happens to be right in the middle. If you average 10 and 12, you get 11. So in this case, the simple answer should be 50% Brazilian, 50% Turkish. Or a 1 to 1 ratio. If you want a total of 1 kg mixture, than .5 kilograms each.

OpenStudy (saifoo.khan):

are u still stuck?

OpenStudy (anonymous):

yest s

OpenStudy (saifoo.khan):

You can start this problem thinking in terms of K amount of Kenyan coffee @ $10 per kilogram + V amount of Venezuelan coffee @ $12 per kilogram = 360 kg @ $11/kg . The problem asks how many kg of Kenyan coffee in the mixture. Well, it can be said the amount is 360 - V. That is, the amount of Kenyan = known total - unknown amount of Venezuelan coffee. . At first, this may not seem like much help. But in fact, that is one of the key ways to reduce the number of unknown variables in algebra word problems. . K = 360 - V V = V . Substituting in the total cost equation: . 10*(360-V) + 12*(V) = 360*11 . Multiplying through: 3600-10V + 12V = 3960 . Collecting like terms 3600 + 2V = 3960 . Subtracting 3600 from both sides 2V = 360 . Dividing both sides by 2 V = 180 . So, V = 180, which means K = 360-180 = 180. . We have equal amount of K and V in the total, which makes sense because we want the mixture to cost $11/kg and the two components cost $10/kg and $12/kg. $11/kg is the average. . We can check our work by substitution back into the equation to see if our result produces the given result. 180*10 + 180*12 = 36*11 1800 + 2160 = 3960 = 36*11 . Done.